OpenStudy (anonymous):

Five percent of what number is 2.5?

OpenStudy (anonymous):

I WILL BECOME A FAN AND GIVE MEDALS

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Make an equation. 5% of what number is 2.5 '5%' is the same as 0.05 'of' indicates multiplication 'what number' indicates an unknown variable, 'x' 'is' indicates an equal sign. So we have: 0.05 * x = 2.5 0.05x = 2.5 Divide 0.05 to both sides, what's 2.5 / 0.05?

OpenStudy (anonymous):

Let 5% of x = 2.5 \[\frac5 {100} \times x = 2.5\] \[x = \frac{2.5 \times 100}{5} = 50\]

OpenStudy (anonymous):

Thirty percent of what number is 13.5?

OpenStudy (anonymous):

Let 30 % of x = 13.5 \[\frac {30}{100} \times x = 13.5\]\[x= \frac{13.5 \times 100}{30} = ? \]

OpenStudy (anonymous):

x = 45